Look up the latest financial statements for a non-financial company of your choice and calculate the following ratios for the latest year a. Return on capital b. Return on equity c. Operating profit margin. d. Days in inventory. e. Debt ratio. f. Times-interest-earned. g. Current ratio h. Quick ratio. Discuss your results.

STATEMENT OF CHANGES IN EQUITY Company Group 30/6/2018 30/6/2017 30/6/2018 30/6/2017 Equity balance at the years beginning (1.1.2018 &1.1.2017 196.522.65 214.497,15 247.940,31 (39.279,40) (13.754,12) 20.879,91 4.605,88 244.271,05 (39.279,40) (2856684) (11.815,84) (19.087,28) (28.566,84) Dividends paid Profit / (Loss) Other total comprehensive income after tax IFRS 9, 15 impact Balance at the end of period (30.06.2018 & 30.06.2017 20.879,91 14.608,74 180.916,06 (20.037,54) (41.548,63) 0,00 41.362,24 0,00 125.480,80 230.392,59 54.118,07 ADDITIONAL DATA & INFORMATION

STATEMENT OF COMPREHENSIVE INCOME Company Group 01/01 455.713,71 (23.948,63) 01/01 01/01-3006/2018 01/01-3006/2017 30/06/2018 30v06/2017 Revenue Gross protit EBIT Profit(loss) before tax Income tax Profit /(loss) after tax (a) Other Total Comprehensive Income (b) Total Comprehensive Income (a)+(b) 368.460,71 (10.744,31) (21.600,96 15.825,13) 371.435,91 450.674,50 4.612.15 (4.120,12) (18.691,45) (25.404,61) 6.317,34 19.087,28) 41.362,24 (1.400,93) (19.421,20) (18.257,23 (26.691,08) 4.009,29 11.815,84 20.879,91 9.064,07 4.503.11 6.653,54 13.754,12 20.879,91 20.037 41.548,62 61.586,15 449,52 7.125,7 Basic earnings after tax per share in EBITDA (0,1926) (14.742,13) 0,2806) 13) (10.453,68) 8.831,99 (10.225,23)

Inv Purchases of assets Sales of tangible & intangible assets Downpayments for purchases of tangible assets Purchace of financial assets Sale of financial assets Interest and other financial income received Net cash flows from investing activities (b) 7,87 (33.583,61) (5.654,50) 0.00 398,01 (3.799,65) 1.570,19) 0,00 0,00 (450,00) 598,49 219,41 3.820,64) (1.575,55) 0,00 0,00 (450,00) 0,00 .322,57 238,95 7,87 (33.583,61) (5.654,50) 426,38 Dividends paid nges in financial leases capital Net cash flows from financing activities (c) Net increase/(decrease) in cash and cash equivalents (35.474,21) (28.523,05) 4.844,62) (5.353,63) (40.318,83) (33.876,68) 86.696,25 82.712,14 (35.474,21) (28.523,05) (4.844,62) (5.353,63) 33.876,68 80.212,62 93.091,47 40.318,83 h and cash equivalents at the beginning of the period 222.403,88 194.454,36 300.931,52 248.477,75 1.980,29 (7.968,94) 383.124,42 333.600,28 Net foreign exchange differences 1.960,42 (7.664,04) h and cash equivalents at the end of the period 311.060,56 269.502,46

Company Group 30/6/2018 30V6/2017 30/6/2018 30/6/2017 Profit/ (loss) before taxes Plus less) adjustments for: (15.825,13) (25.404,61) (18.257,23) (26.691,08) 8.466,21 (793, 1.169.25 (6.111,98) 5.404,80 747,28) (565,27) 1.873,64 8.831,99 Foreign exchange differences (Profit) loss from investing activities Finance Cost 9.206,49 8.967,52 (1.372,85) .184,64 (6.097,97) 5.985,42 (775,65) (712,34) 1114,60 .902,33 1.083,42 h flows from operating activities before changes in capital (Increase) /Decrease in inventories (Increase) Decrease in trade & other receivables Increase/(Decrease) in payables (other than banks) Interest expenses paid ncome tax payments Net cash flows from operating activities (a) 1.139,64)(519,04) (1.357,09) 835,22 (58.283,20) (63.406,78) (59.797,86) (69.986,53) 241.453,28 211.692,96 (959,77) 1.238,14) 243.560,49 196.490,18 (928,59) 1.209,45) 0,00 507,8 0,00 507,8 169.646,97 117.791,11 163.155,94 127.432,18

FINANCIAL POSITION Company Group 30/6/2018 31/12/2017 30/6/2018 31/12/2017 ASSETS Tangible assets Investments in subsidaries Goodwill Intagible assets Other non current assets Inventories Customers and other trade receivables 94.725,87 94.326,38 72.416,56 72.416,56 0,00 27.234,38 27.863,32 62.834,52 21.988,14 8.964,05 43.637,42 103.522,29 09.945,45 215.195,95 77.085,21 50.153,65 797.983,09 594.430,34 94.510,24 94.191,86 0,00 39.756,30 39.756,30 44.212,46 45.113,82 67.378,62 25.896,04 16.550,34 5.193,25 57 623,38 115.556,86 382.009,31 293.723,58 78.646,11 52.149,87 880.686,77 681.581,59 0,00 0,00 10.103,69 Cash and cash equivalents Other current assets TOTAL ASSETS Share capital Additional paid-in capital and reserves Total shareholders equity (a) Long term bank loans 46.421,11 46.421,11 34.494,95 150.101,54 180.916,06 196.522,65 0,00 46.421,11 46.421,1 83.971,48 201.519,19 230.392,59 247.940,31 0,00 0,00 0,00 Provisions and other long term liabilities Short term bank loans Other short term liabilities 69.517,09 64.497,93 0,00 547.549,95 333.409,76 617.067,04 397.907,70 76.117,02 71.439,48 0,00 574.177,15 362.201,80 650.294,18 433.641,28 0,00 0,00 Total liabilities (b) EQUITY AND LIABILITIES (c) ((b) 797.983,09 594.430,34 880.686,77 681.581,59
Answer #1

Calculation of ratios :

1. Return on capital :

Return on capital employed (ROCE) is a financial ratio that measures a company's profitability and the efficiency with which its capital is employed. ROCE is calculated as:

ROCE = Earnings Before Interest and Tax (EBIT) / Capital Employed

= 2160096 / 25043314

= 8.6254%

Result: It is higher the better.

2. Return on equity :

Return on equity measures a company's profit as a percentage of the combined total worth of all ownership interests in the company

=Net income / shareholders equity

= 1181584/18091606

= 6.5311%

Result : Higher the better

3. Operating profit margin

Operating Profit Margin is a profitability or performance ratio used to calculate the percentage of profit a company produces from its operations, prior to subtracting taxes and interest charges. It is calculated by dividing

=operating profit / total revenue

and expressing as a percentage. The margin is also known as EBIT (Earnings Before Interest and Tax) Margin.

=2160096 /36846071

=5.8624%

Result : Higher the better.

4. Days in inventory ratio:

Days in inventory (also known as 'Inventory Days of Supply(DoS)', 'Days Inventory Outstanding' or 'the Inventory Period') is an efficiency ratio that measures the average number of days the company holds its inventory before selling it. The ratio measures the number of days funds are tied up in inventory. Inventory levels (measured at cost) are divided by sales per day (also measured at cost rather than selling price.)

The formula for DII is:

{\displaystyle DII={\dfrac {average~inventory}{COGS/Days}}}

where the average inventory is the average of inventory levels at the beginning and end of an accounting period, and COGS/day is calculated by dividing the total cost of goods sold per year by 365 days.

Avg inventory = opening inventory + closing inventory / 2

= 891005+1010369/2

= 950687

COGS/day = 891005+0 - 1010969= - 119964/365 = - 328 day

Result : The ratio is adverse, as it is negative.

5. Debt ratio :

The debt ratio is a financial ratio that measures the extent of a company’s leverage. The debt ratio is defined as the ratio of total debt to total assets, expressed as a decimal or percentage. It can be interpreted as the proportion of a company’s assets that are financed by debt. formula for calculating the debt ratio is:

= total liabilities / total assets

= 61706704/ 79,798,309

= 77.328%

Result : financial leverage is high as it is around 77%

6. Times interest earned :

Times interest earned (TIE) is a metric used to measure a company's ability to meet its debt obligations. The formula is calculated by taking a company's earnings before interest and taxes (EBIT) and dividing it by the total interest payable on bonds and other contractual debt. TIE indicates how many times a company can cover its interest charges on a pretax earnings basis.

= EBIT / Total interest payable

= 2160096 /92859

= 23.26

Result : ratio is adequate  

7. Current ratio:

The current ratio is a popular metric used all across the industry to assess a company's short-term liquidity with respect to its available assets and the pending liabilities.

Current assets include cash, cash equivalents, accounts receivable, stock inventory, marketable securities, pre-paid liabilities and other liquid assets

On the balance sheet, current liabilities are typically presented as follows: the principal portion of notes payable due within one year, accounts payable, and then other current liabilities, such as income taxes payable, interest payable, et al.

The formula is current assets / current liabilities

= 54077177/ 54754995

= 0.987

Result : The ideal current ratio is 2: 1. Our ratio is lower than ideal.

8. Quick ratio :

The quick ratio or acid test ratio is a liquidity ratio that measures the ability of a company to pay its current liabilities when they come due with only quick assets. Quick assets are current assets that can be converted to cash within 90 days or in the short-term. Cash, cash equivalents, short-term investments or marketable securities, and current accounts receivable are considered quick assets.

= quick assets /current liabilities

= 54,077,177 /54754995

= 0.987

Result : same as current ratio.